China had $40.00 to spend on herself.  She spent 1/5 of the money on a chicken tender meal, 1/4 for a set of earrings, and 1/2 of it on a bracelet.       How much money does she have left over?

1
Expert's answer
2020-12-21T18:00:22-0500

China spent "\frac{1}{5}+\frac{1}{4}+\frac{1}{2}=\frac{4}{20}+\frac{5}{20}+\frac{10}{20}=\frac{4+5+10}{20}=\frac{19}{20}" of money chicken tender meal , set of earrings

, bracelet and "1-\frac{19}{20}=\frac{20-19}{20}=\frac{1}{20}" of money china has left over.

It is "\frac{1}{20}\times \$ 40 =\$2".


Answer: $2